OpenStudy (anonymous):

If tan x= -5/12 and x is in quadrant 2, then sin 2x = ?

OpenStudy (anonymous):

3/5

OpenStudy (anonymous):

2(-1/√10)(-3/√10) = 6/10 = 3/5

OpenStudy (anonymous):

Do u understand?

OpenStudy (anonymous):

no not at all..